OpenStudy (anonymous):

The label on a juice carton says there are 50 calories in a 4-ounce serving of juice. If Arthur drinks 12 ounces of juice, how many calories will he consume? A. 48 calories B. 58 calories C. 150 calories D. 200 calories

OpenStudy (micahm):

www.fda.gov/.../GuidanceDocumentsReg...

OpenStudy (anonymous):

c

OpenStudy (anonymous):

thank you @cooolkilla02!

OpenStudy (anonymous):

12/4=3 Servings 3*50=150 calories

OpenStudy (anonymous):

is that right though?

OpenStudy (anonymous):

If 4 oz= 50 what is 12oz= so whats 12/4?

OpenStudy (anonymous):

Yes @DoctorWholover23

OpenStudy (anonymous):

can i have a medal @DoctorWholover23

OpenStudy (anonymous):

okay thank you! and yes you can!

OpenStudy (anonymous):

3 right? so then you multiply 3 times the calories, which is 50

OpenStudy (anonymous):

ok last step 3*50=

OpenStudy (anonymous):

ok thank you @EveTheAwsome and 3*50=150

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

:)